OpenStudy (anonymous):

If the initial velocity of the car in the previous problem (the car accelerates at 5.00 km/hr/s for 3.50 sec) is 22.6 km/hr, what is the final velocity of the car?

OpenStudy (anonymous):

Let final velocity be x Acceleration= final velocity - initial velocity/time 5=x-22.6/3.5 5*3.5=x-22.6 17.5+22.6=x 40.10=x Therefore, final velocity is 40.10 km/hr